What is the magnitude of the Electric force 55 × 10 uC charge exerts on a 8.5 × 10^3 C charge located 0.75 m away ?


1
Expert's answer
2021-06-09T08:05:38-0400

The magnitude of the Electric force is given by the Coulomb's law:


"F = k\dfrac{q_1q_2}{r^2}"

where "q_1 = 55\times 10uC = 550\times 10^{-6}C, q_2 = 8.5\times 10^3C, r = 0.75m" and "k = 9\times 10^9N\cdot m^2/C^2". Thus, obtain:


"F = 9\times 10^9\cdot \dfrac{550\times 10^{-6}\cdot 8.5\times 10^{3}}{0.75^2}= 7.48\times 10^{10}N"

Answer. "7.48\times 10^{10}N".